# Ab initio uncertainty quantification in scattering analysis of microscopy

### v2

@@ VI.1 Materials and Methods
The sample is imaged using a Zeiss Axio Observer 7 microscope in fluorescence mode using a Colibri 7 light source, and standard GFP filter sets. The images are captured with a 20x objective, with a numerical aperture of 0.8. A typical image size is 512 $\times$ 512 pixels with a pixel size of 0.29 $\mu$m/pixel and T = 500 time steps with a step size of 0.0309 seconds.
@@ VI.2 Diffusive motion with different particle size and number density
The MSD of a diffusive process can be related to viscosity by the Stokes-Einstein equation [74]:
@@
where $k_{B}=1.38\times 10^{-23}J/K$ is the Boltzmann constant, $T_{a}$ is the absolute temperature, $r_{p}$ is the radius of the particle. Keeping $T_{a}$, $\eta$ constant, the slope of $\langle\Delta x^{2}(\Delta t)\rangle$ versus $\Delta t$ decreases proportionally to increasing particle radius $r_{p}$. We prepare several of the same composition but mix them with $2r_{p}=$ 1 $\mu$m, 200 nm, and 100 nm particles, separately. All particles are used at volume fraction $\phi$ = 1×$\times 10^{-4}$, but the number densities of the 200 nm and 100 nm particles are much higher, as shown in the insets in Fig. 6.
@@
We compare MPT, AIUQ approaches and two fitting approaches based on image structure functions. For AIUQ approaches, we show the performance based on all q or reduced q to accelerate the computation. For all real experiments in Sec. VI.2-Sec. VI.4, the smallest $J_{0}$ such that $\sum^{J_{0}}_{j=1}A_{j}/\sum^{J}_{j=1}A_{j}\geq 1-\varepsilon$ with $\varepsilon=0.99$ is used as a default choice of the wavevectors for reducing the computational cost in the AIUQ with reduced q approach. Such choice leads to $J_{0}\approx 0.25J$, which reduces the computational cost by more than 10 times compared to AIUQ on all $q$, suitable for scalably computing a large number of experimental data, especially for those discussed in Sec. VI.3. For the three sets of experiments discussed here, we assume an FBM to parameterize ISF in AIUQ approaches and fitting image structure function approaches instead of BM to test whether the model can identify the Newtonian behavior of the fluids, as well as validate that other potential factors, such as drifts, do not have a large impact on the results.
@@
As shown in Fig. 6, parameter estimation by fitting $D(q,\Delta t)$ is less accurate than the by AIUQ approaches, based on either all $q$ or reduced $q$, compared to the reference value (black curves). This result highlights that the wavevector range needs to be carefully selected when fitting image structure function with conventional DDM, but this problem is resolved using AIUQ, which uses the maximum marginal likelihood estimator. For nanoparticles with $2r_{p}=$ 100 and 200 nm in panels (b)-(c), which are below the diffraction limit, MPT systematically underestimates the particle movements compared to the truth, due to higher concentrations and faster dynamics. MPT computes the ensemble MSD by linking the trajectories and averaging the displacements from all particles. When dealing with an optically dense sample, the probability of erroneously linking two nearby particles increases. Furthermore, the progressively worsening performance of MPT is attributed to averaging fewer particle steps towards the larger $\Delta t$’s. We note that quite often, MPT only utilizes the first 10-20% of the $\Delta t$’s, requiring substantially larger time range to be measured.
@@
The estimated MSDs by AIUQ approaches are closer to the reference values than MPT for scenarios in Fig. 6(b)-(c), as AIUQ overcomes the difficulty of separating and linking a large number of particles. Optically dense systems are not uncommon in experiments. In some experiments, for instance, larger particles are incompatible with the system, when the density and viscosity of the continuous phase are both low - probe particles tend to sediment unless they are small enough to be dispersed by Brownian forces [44]. The second scenario is when higher moduli are expected and thus using small probe particles is necessary to produce detectable displacements. A third scenario is when the system already contains probes, such as tracking phase-separated regions inside organelles [75], where the size of the particles cannot be adjusted. AIUQ approaches are well-suited for analyzing optically dense samples without the need for specifying wavevector ranges or tuning parameters.
@@
Finally, all approaches seem to slightly overestimate MSD at a large $\Delta t$ for the experiment with $2r_{p}$ = 1 $\mu m$ particles shown in Fig. 6(a). This is because small drift disproportionally impacts the larger probes, which move more slowly under the thermal fluctuation. Using smaller particles such as those in Fig. 6(b)-(c) can mitigate these impacts. Another way is to model the drift from ISF and integrate it in AIUQ estimation.
@@ VI.3 Automated estimation of gelling point of a perfect network
Figure 7: (a) Schematics showing the reaction between tetraPEG-SG and tetraPEG-NH2. (b) Multiple particle tracking of the time series ensemble mean-squared displacement of the probes embedded in the mixed tetra-functional PEG plotted against lag time $\Delta t$. (c) AIUQ analysis of the same data as (b), denoted by the same color but with dotted lines. (d) Superposition of the data in (b) using MPT data. (e) Shift factors $a$ (red) and $b$ (blue) for pregel (circles), and postgel (triangles). (f) Estimated $\rho$ parameters in the OU process from a sequence of videos at different times denoted by black circles. The black solid line denotes the fit from a generalized logistic curve where $R^{2}_{adjusted}>0.99$.
@@
The shifting is carried out by multiplying the MSD curves and the $\Delta t$ by coefficients $a$ and $b$, determined by visual inspection. Once done, the MSD curves are collapsed into two distinct branches, one pre-gel and one post-gel. Each dataset (for a particular gelling time) should exhibit some overlap with the subsequent one. The MSD at the gelling point is a critical curve between these two branches, which exhibits power-law-like behavior. The critical gelling exponent is defined as $\alpha=\frac{d\ln\langle\Delta x^{2}(\Delta t)\rangle}{d\ln\Delta t}$ [9, 77], where $\alpha$ is constant over all $\Delta t$, identical to the FBM model. Different systems experience gelation with distinct critical power-law exponents, as noted in [84]. These exponents can span from as high as $\alpha=0.88$ [85] to as low as $\alpha=0.16$ [82].
A recent approach [86] aimed to automate the process, using MSD versus $\Delta t$ curves estimated by MPT. However, these MSD curves are non-smooth and therefore extra efforts are required to smooth the MSD curves before implementing the superposition. DDM has also been utilized to explore gelation behaviors, though manually superposing the MSD curves is still needed. In [78], DDM is used as an initial screening step and the authors identify gelation when displacement falls below a threshold, at which point it is challenging to analyze subtle displacements using DDM, whereas the pre-specified wavevector range can affect the estimation. In [28], the authors conducted MSD fitting using both MPT and DDM. Then they extracted the log-slope of MSD curves from MPT and subsequently fit a logit function to these slopes to identify gelation.
@@
Next, we demonstrate using AIUQ to automate the gelling point determination. The viscoelastic solid can be modeled by an OU process, with $\mbox{MSD}(\Delta t)=\sigma^{2}_{OU}(1-\rho^{\Delta t})$, which can capture the plateau and the reducing gradient of the MSD curve at large $\Delta t$’s. As the number of experiments is large, we use AIUQ with reduced $q$ for estimating the ISF. The estimated MSDs from the AIUQ approach with an OU model are shown in Fig. 7(c), which resemble curves from MPT curves from Fig. 7(b). The two parameters from the OU models, $\rho$ and $\sigma^{2}_{OU}$, determine the shape of the MSDs. These estimated parameters can be further processed to deduce the gel point of the material. In Fig. 7(e), we plot the estimated parameter $\rho$ from the OU model for each experiment. The estimated $\rho$ gradually decreases, due to the caging effect from the network that traps particles at longer $\Delta t$’s. Before gelling, the absolute change of estimated $\rho$ increases, and the absolute change decreases after gelling. The rapid drop of $\rho$ from 1 to 0 indicates the sol-to-gel transition. To find the gelling point, we fit a generalized logistic function, $\rho(t)=\exp(-c_{1}(t-c_{2}))/(1+\exp(-c_{1}(t-c_{2})))$, where $c_{1}$ and $c_{2}$ are determined by minimizing a $L_{1}$ loss with respect to $\rho$. The $L_{1}$ loss is more robust than the $L_{2}$ loss, and here both give similar estimations. The fit is shown by the black solid line, which characterizes the change of $\rho$, and defines the gelling time to be the time point with the largest intermediate change in $\rho$. The inflection point is found to be 43.4 minutes using this fit, which is similar to the estimate $T_{gel}=44$ min in the previous study [87].
@@
To extract the critical gelling exponent, we again compare two approaches. First, we plot the log shift factors $\log a$, $\log b$ against $\log$ of the extent of reaction $c$, defined as $c=\frac{|{T-T_{gel}}|}{T_{gel}}$, where $T_{gel}$ is the gelling time, hence $c$ determines distance to the gelling point [88]. Given $a\propto c^{-\alpha_{a}}$, and $b\propto c^{-\alpha_{b}}$, the scaling exponent is the ratio of these two exponents $\alpha=\alpha_{a}/\alpha_{b}$. This way, we obtain the scaling exponent for the 4-arm PEG-NH2 and 4-arm PEG-SG system to be $\alpha$ = 0.739. The slope is plotted in Fig. 7(d) for reference. A second way is similar to what has been described in [28], the authors fit a logistic function to the relaxation exponent of the MSD. The same information can be obtained by fitting the fractional Brownian motion to all of the MSD curves, and we found that the critical exponent of $\alpha$ = 0.74 at gel point t = 43.4 min. Thus, the scattering analysis of microscopy by the AIUQ approach can be used to automatically extract the critical quantities of gelling systems.
@@
The AIUQ approach extends the boundary of previous techniques in the post-gel branch of the MSD for systems undergoing gelation. Compared to MPT and conventional curve shifting for superposition by hand, AIUQ lifts the hurdles of strenuous analysis for a large number of videos by providing an automated estimation of gelling time, without the need of specifying wavevector range or superposing the curves. Thus the new tool can be deployed for automatically estimating the gelation time by a sequences of microscopy videos, at a given formulation and experiment condition (e.g. temperature, pH, etc). Along with high-throughput experiments, this data-processing technique can be integrated with Bayesian optimization and active learning approaches [89, 90] to optimize the compositions or material designs to achieve ideal gelling time and properties.
@@ VI.4 Anisotropic diffusion in lyotropic LC
[…]
@@
[…]
An example of the anisotropic particle trajectory is shown in Fig. 8(d) inset. A probe particle moves predominantly along the x direction (sky blue) compared to y (purple). The previous MSD by particle tracking results from previous work [91] is plotted by triangle symbols in Fig. 8. The authors also determined the viscosities parallel and perpendicular to the alignment direction, which are plotted as solid and dashed lines in Fig. 8. Since the log-slope $\alpha=\frac{d\ln\langle\Delta x^{2}(\Delta t)\rangle}{d\ln\Delta t}<1$, we use the FBM model in the AIUQ approach.
MSDs by MPT and AIUQ are shown in the lin-lin plot in Fig. 8(d) and then again in log-log in Fig. 8(e). As the computation cost for anisotropic processes is higher than isotropic processes, due to the larger number of parameters, and distinct ISFs at each wavevectors, we explore this kind of behavior with a reduced number of $q$’s, but the estimation from AIUQ with all $q$ is similar. Both MPT and AIUQ effectively capture large distinctions in diffusion along with direction, and the results are similar to the values reported in [91], as shown in Fig. 8(d).
When examining the log-log plot of MSD, the MSD by MPT suggests a subdiffusive region at a small $\Delta t$. In a similar system, a subdiffusive-to-diffusive transition was observed at the time scale of twist relaxation, which is on the order of $\Delta t\approx$ 100 s [93], but with much larger particles ($2a\approx$ 7 $\mu m$). The authors postulate that subdiffusive behavior can be attributed to the restoring forces on the particle from incurred elastic free energy costs, partially offsetting the displacement from thermally driven fluctuations. Given that the model used in AIUQ for our case is FBM, we constrain the log-log plot of the MSD to be linear. As noted in [91], physical models that can explain anistropic subdiffusive behaviors in nematic liquid crystal has yet to be constructed. This indicates that a model-free or nonparameteric of ISF is appealing when the physical model of ISF is unavailable, whereas these approaches typically require a specified wavevector range for inverting the image structure function in DDM [26, 45, 44]. Whereas utilizing the plateau or variability of the image structure curves can be helpful for identifying the wavevector range, it is of interest to integrate the model-free ISF in AIUQ approaches for inverse estimation.

### v4

@@ VI.1 Materials and Methods
The samples are imaged using a Zeiss Axio Observer 7 microscope in fluorescence mode using a Colibri 7 light source, and standard GFP filter sets. The images are captured with a 20x objective, with a numerical aperture of 0.8. A typical image size is 512 $\times$ 512 pixels with a pixel size of 0.29 $\mu$m/pixel and n = 500 time steps with a step size of 0.0309 seconds.
We have simulated 2D motion for simplicity, but the probe movements are intrinsically 3D in real experiments. For these analyses, the motion of the particle is projected onto the plane of the focus and only 2D trajectories are obtained and analyzed [76]. There are cases for which this simplification does not work and the observed dynamics are affected by axial motion, such as shown in [5]. Nonetheless, in all the cases shown here, the axial displacement is either negligible, or the motion is isotropic in all three dimensions so a 2D treatment of the 3D scenario is warranted.
@@ VI.2 Diffusive motion with different particle size and number density
The MSD of a diffusive process can be related to viscosity $\eta$ by the Stokes-Einstein equation [77]:
@@
where $k_{B}=1.38\times 10^{-23}J/K$ is the Boltzmann constant, $T_{a}$ is the absolute temperature, $r_{p}$ is the radius of the particle. Keeping $T_{a}$, $\eta$ constant, the slope of $\langle\Delta x^{2}(\Delta t)\rangle$ versus $\Delta t$ decreases proportionally to increasing particle radius $r_{p}$. We prepare several samples of the same composition but adding particles of different sizes $2r_{p}=$ 1 $\mu$m, 200 nm, and 100 nm in each one. All fluorescent image sequences were captured at particle volume fraction $\phi$ = 1×$\times 10^{-4}$, but the number densities of the 200 nm and 100 nm particles are much higher, as shown in the insets in Fig. 6. The bright field image sequence was captured at particle volume fraction $\phi$ = 2×$\times 10^{-3}$.
@@
We compare MPT, AIUQ approaches and two fitting approaches based on image structure functions. For AIUQ approaches, we show the performance based on all q or reduced q to accelerate the computation. For all real experiments in Sec. VI.2-Sec. VI.4, the smallest $J_{0}$ such that $\sum^{J_{0}}_{j=1}A_{j}/\sum^{J}_{j=1}A_{j}\geq 1-\varepsilon$ with $\varepsilon=0.99$ is used as a default choice of the wave vectors for reducing the computational cost in the AIUQ with reduced q approach. Such choice leads to $J_{0}\approx 0.25J$, which reduces the computational cost by more than 10 times compared to AIUQ on all $q$, suitable for scalably computing a large number of experimental data, especially for those discussed in Sec. VI.3. For the three sets of experiments discussed here, we assume the FBM model to parameterize ISF in AIUQ approaches and fitting image structure function approaches instead of BM to test whether the model can identify the Newtonian behavior of the fluids, as well as validate that other potential factors, such as drifts, do not have a large impact on the results.
@@
As shown in Fig. 6, parameter estimation by fitting $D(q,\Delta t)$ is less accurate than the by AIUQ approaches, based on either all $q$ or reduced $q$, compared to the reference value (black curves). This result highlights that the wave vector range needs to be carefully selected when fitting image structure function with conventional DDM, but this problem is resolved using AIUQ, which uses the maximum marginal likelihood estimator. For nanoparticles with $2r_{p}=$ 100 and 200 nm in panels (b)-(d), which are all below the diffraction limit, MPT systematically underestimates the particle movements compared to the truth, due to higher concentrations and faster dynamics. MPT computes the ensemble MSD by linking the trajectories and averaging the displacements from all particles. When dealing with an optically dense sample, the probability of erroneously linking two nearby particles increases. Furthermore, the progressively worsening performance of MPT is attributed to averaging fewer particle steps towards the larger $\Delta t$’s. We note that quite often, MPT only utilizes the first 10-20% of the $\Delta t$’s, requiring substantially larger time range to be measured. Furthermore, Fourier-based algorithms are better for MPT at analyzing data sets with a low signal-to-noise ratio, such as the brightfield image shown in panel (d), where particles are unresolvable by the eye. As further verifications, estimates for $A(q)$ are presented in Fig. S2, and observed image structure function $D(q,\Delta t)$ and the estimated $D(q,\Delta t)$ using either AIUQ or DDM are shown in Figs. S3 and S4.
@@
The estimated MSDs by AIUQ approaches are closer to the reference values than MPT for scenarios in Fig. 6(b)-(d), as AIUQ overcomes the difficulty of separating and linking a large number of particles. Optically dense systems are not uncommon in experiments. In some experiments, for instance, larger particles are incompatible with the system, when the density and viscosity of the continuous phase are both low - probe particles tend to sediment unless they are small enough to be dispersed by Brownian forces [45]. The second scenario is when higher moduli are expected and thus using small probe particles is necessary to produce detectable displacements. A third scenario is when the system already contains probes, such as tracking phase-separated regions inside organelles [78], where the size of the particles cannot be adjusted. AIUQ approaches are well-suited for analyzing optically dense samples without the need for introducing fluorescent particles, specifying wave-vector ranges, or tuning parameters.
@@
Finally, all approaches seem to slightly overestimate MSD at a large $\Delta t$ for the experiment with $2r_{p}$ = 1 $\mu m$ particles shown in Fig. 6(a). This is because small drift disproportionally impacts the larger probes, which move more slowly under the thermal fluctuation. Using smaller particles such as those in Fig. 6(b)-(c) can mitigate these impacts. Another way is to model the drift from ISF and integrate it into AIUQ estimation.
@@ VI.3 Automated estimation of gelling point of a perfect network
Figure 7: (a) Schematics showing the reaction between tetraPEG-SG and tetraPEG-NH2. (b) Multiple particle tracking of the time series ensemble mean-squared displacement of the probes embedded in the mixed tetra-functional PEG plotted against lag time $\Delta t$ in log-log space. (c) AIUQ analysis of the same data as (b), denoted by the same color but with dotted lines. (d) Superposition of the data in (b) using MPT data. (e) Shift factors $a$ (red) and $b$ (blue) for pregel (circles), and postgel (triangles). (f) Estimated $\rho$ parameters in the OU process from a sequence of videos at different times denoted by black circles. The black solid line denotes the fit from a generalized logistic curve where $R^{2}_{adjusted}>0.99$.
@@
The shifting is carried out by multiplying the MSD curves and the $\Delta t$ by coefficients $a$ and $b$, determined by visual inspection. Once done, the MSD curves are collapsed into two distinct branches, one pre-gel and one post-gel. Each dataset (for a particular gelling time) should exhibit some overlap with the subsequent one. The MSD at the gelling point is a critical curve between these two branches, which exhibits power-law-like behavior. The critical gelling exponent is defined as $\alpha=\frac{d\log\langle\Delta x^{2}(\Delta t)\rangle}{d\log\Delta t}$ [9, 80], where $\alpha$ is constant over all $\Delta t$, identical to the FBM model. Different systems experience gelation with distinct critical power-law exponents, as noted in [87]. These exponents can span from as high as $\alpha=0.88$ [88] to as low as $\alpha=0.16$ [85].
A recent approach [89] aimed to automate the process, using MSD versus $\Delta t$ curves estimated by MPT. However, these MSD curves are non-smooth and therefore extra efforts are required to smooth the MSD curves before implementing the superposition. DDM has also been utilized to explore gelation behaviors, though manually superposing the MSD curves is still needed. In [81], DDM is used as an initial screening step and the authors identify gelation when displacement falls below a threshold, at which point it is challenging to analyze subtle displacements using DDM, whereas the pre-specified wave-vector range can affect the estimation. In [29], the authors conducted MSD fitting using both MPT and DDM. Then they extracted the log-slope of MSD curves from MPT and subsequently fit a logit function to these slopes to identify gelation.
@@
Next, we demonstrate using AIUQ to automate the gelling point determination. The viscoelastic solid can be modeled by an OU process, with $\mbox{MSD}(\Delta t)=\sigma^{2}_{OU}(1-\rho^{\Delta t})$, which can capture the plateau and the reducing gradient of the MSD curve at large $\Delta t$’s. As the number of experiments is large, we use AIUQ with reduced $q$ for estimating the ISF. The estimated MSDs from the AIUQ approach with an OU model are shown in Fig. 7(c), which resemble curves from MPT curves from Fig. 7(b). The two parameters from the OU models, $\rho$ and $\sigma^{2}_{OU}$, determine the shape of the MSDs. These estimated parameters can be further processed to deduce the gel point of the material. In Fig. 7(e), we plot the estimated parameter $\rho$ from the OU model for each experiment. The estimated $\rho$ gradually decreases, due to the caging effect from the network that traps particles at longer $\Delta t$’s. Before gelling, the absolute change of estimated $\rho$ increases, and the absolute change decreases after gelling. The rapid drop of $\rho$ from 1 to 0 indicates the sol-to-gel transition. To find the gelling point, we fit a generalized logistic function, $\rho(t)=\exp(-c_{1}(t-c_{2}))/(1+\exp(-c_{1}(t-c_{2})))$, where $t$ is the reaction time, $c_{1}$ and $c_{2}$ are determined by minimizing a $L_{1}$ loss with respect to $\rho$. The $L_{1}$ loss is more robust than the $L_{2}$ loss, and here both give similar estimations. The fit is shown by the black solid line, which characterizes the change of $\rho$, and defines the gelling time to be the time point with the largest intermediate change in $\rho$. The inflection point is found to be $43.4$ minutes using this fit, which is similar to the estimate $44$ minutes in the previous study [90].
@@
To extract the critical gelling exponent, we again compare two approaches. For the MPT approach, we plot the log shift factors $\log a$, $\log b$ against $\log$ of the extent of reaction $c$, defined as $c=\frac{|{t-t_{gel}}|}{t_{gel}}$, where $t_{gel}$ is the estimated gelling time, hence $c$ determines distance to the gelling point [91]. Given $a\propto c^{-\alpha_{a}}$, and $b\propto c^{-\alpha_{b}}$, the scaling exponent is the ratio of these two exponents $\alpha=\alpha_{a}/\alpha_{b}$. This way, we obtain the scaling exponent for the 4-arm PEG-NH2 and 4-arm PEG-SG system to be $\alpha$ = 0.739. The slope is plotted in Fig. 7(d) for reference. A second way is similar to what has been described in [29], the authors fit a logistic function to the relaxation exponent of the MSD. The same information can be obtained in AIUQ by fitting the fractional Brownian motion to the MSD curves closest to the gelling point, and we found that the critical exponent of $\alpha$ = 0.74 for the estimated gelling point t = 43.4 min. Thus, the scattering analysis of microscopy by the AIUQ approach can be used to automatically extract the critical quantities of gelling systems.
@@
The AIUQ approach extends the boundary of previous techniques in the post-gel branch of the MSD for systems undergoing gelation. Compared to MPT and conventional curve shifting for superposition by hand, AIUQ lifts the hurdles of strenuous analysis for a large number of videos by providing an automated estimation of gelling time, without the need of specifying wave-vector range or superposing the curves. Thus the new tool can be deployed for automatically estimating the gelation time by a sequences of microscopy videos, at a given formulation and experiment condition (e.g. temperature, pH, etc). Along with high-throughput experiments, this data-processing technique can be integrated with Bayesian optimization and active learning approaches [92, 93] to optimize the compositions or material designs to achieve ideal gelling time and properties.
@@ VI.4 Anisotropic diffusion in lyotropic LC
Figure 8: Anisotropic diffusion in DSCG. (a) Schematics of a probe particle moving amongst assembled stacks of DSCG, not drawn to scale. (b-c) Crossed-polarizer images when the channel is aligned either 45o in (b) or parallel to the alignment direction of the channel in (c). The double-sided arrows denote the direction of the polarizer and analyzer. (d-e) Method comparisons between AIUQ fitting (filled symbols) and MPT tracking (solid lines) of the image stack are presented in (d) lin-lin plot and (e) log-log plot. Open circles and triangles denote data reported in [94]. Black solid and dashed lines denote diffusive trends plotted from viscosity values reported from the same paper. The inset in (d) shows an example particle trajectory. The shaded region denotes 95% confidence interval estimated using AIUQ reduced q.
@@
We disperse probes $2r_{p}=$ 200 nm at volume density $\phi$ = 8×$\times 10^{-5}$ in 16 wtDSCG solution. This concentration is chosen as its iso-nematic transition temperature is above room temperature so no isotropic to nematic transition occurred, as the iso-nematic fronts have been found to trap particles and cause aggregation. Given a density of 1.55 g/cm3 for pure DSCG [97], the solution density is 1.088 g/cm3 which is similar to that of polystyrene particles at 1.055 g/cm3 according to the manufacturer, hence these particles are neutrally buoyant. A small amount of surfactant (Triton X-100, 0.015 wt%) is added to prevent aggregation of the probes [96]. The capillary force upon filling is sufficient to ensure good alignment of the DSCG solution [94]. Uniform nematic alignment was obtained this way, as shown by the polarized optical microscopy images in Fig. 8(b)-(c). In these images, regions aligned parallel or perpendicular to the polarizer or analyzer direction (denoted by double-sided arrows) appear dark, while other orientations appear bright.
An example of the anisotropic particle trajectory is shown in Fig. 8(d) inset. A probe particle moves predominantly along the x direction (sky blue) compared to y (purple). The previous MSD by particle tracking results from previous work [94] is plotted by open circle and triangle symbols in Fig. 8. The authors also determined the viscosities parallel and perpendicular to the alignment direction, which are plotted as solid and dashed lines in Fig. 8. Since the log-slope $\alpha=\frac{d\log\langle\Delta x^{2}(\Delta t)\rangle}{d\log\Delta t}<1$, we use the FBM model in the AIUQ approach.
MSDs by MPT and AIUQ are shown in the lin-lin plot in Fig. 8(d) and then again in log-log in Fig. 8(e). As the computation cost for anisotropic processes is higher than isotropic processes, due to the larger number of parameters, and distinct ISFs at each wave vector, we explore this kind of behavior with a reduced number of $q$’s, but the estimation from AIUQ with all $q$ is similar. Both MPT and AIUQ effectively capture large distinctions in diffusion along with direction, and the results are similar to the values reported in [94], as shown in Fig. 8(d).
When examining the log-log plot of MSD, the MSD by MPT suggests a subdiffusive region at a small $\Delta t$. In a similar system, a subdiffusive-to-diffusive transition was observed at the time scale of twist relaxation, which is on the order of $\Delta t\approx$ 100 s [96], but with much larger particles ($2a\approx$ 7 $\mu m$). The authors postulate that subdiffusive behavior can be attributed to the restoring forces on the particle from incurred elastic free energy costs, partially offsetting the displacement from thermally driven fluctuations. Given that the model used in AIUQ for our case is FBM, we constrain the log-log plot of the MSD to be linear. As noted in [94], physical models that can explain anistropic subdiffusive behaviors in nematic liquid crystal has yet to be constructed. This indicates that a model-free or nonparameteric of ISF is appealing when the physical model of ISF is unavailable, whereas these approaches typically require a specified wave-vector range for inverting the image structure function in DDM [27, 46, 45]. Whereas utilizing the plateau or variability of the image structure curves can be helpful for identifying the wave-vector range, it is of interest to integrate the model-free ISF in AIUQ approaches for inverse estimation.